OpenStudy (anonymous):

Jim opened a new savings account. He deposited $12,000 at 12 percent compounded semiannually. At the start of the fourth year, Jim deposits an additional $50,000 that is also compounded semiannually at 12 percent. At the end of 6 years, what is the balance in Jim's account?

OpenStudy (anonymous):

12,000(1+.12)+50,000(1+.12)^12=208,238.80 ? thats probably not even the right formula...

OpenStudy (dumbcow):

B = P(1+r/n)^nt , Formula for compound interest P = 12,000 and 50,000 r = 0.12 n=2 t = 6 and 3 Total = 12,000(1.06)^12 + 50,000(1.06)^6

OpenStudy (anonymous):

oh. yea i was way off. B=95,072.31

OpenStudy (anonymous):

thank you, the formula was a lot of help.